OpenStudy (anonymous):

What is the value of 12C5 ? A. 95,040 B. 60 C. 792 D. 120

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=12+choose+5

OpenStudy (anonymous):

C

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ k\end{matrix}\right) = \dfrac{n!}{k!(n-k)!}\]

OpenStudy (anonymous):

\[_{12}C_5=\frac{\overbrace{12\times 11\times 10\times 9\times 8}^{\text{ five terms }}}{5\times 4\times 3\times 2}\] cancel first, multiply last, or use wolfram